A disc jockey must play 14 commercial spots during 1 hour of each radio shows. Each commercial is either 30 seconds or 60 seconds long. If the total commercial time during 1 hour is 11 minutes, how many 30 second commercials played that hour? How many 60 second commercials?
Answer by rwm(914)

x+y=14
.5x+y=11
x=6 y=8
check
3+8=11
11=11